### v1

# A Cost-Benefit Analysis of On-Premise Large Language Model Deployment: Breaking Even with Commercial LLM Services

Abstract
Large language models (LLMs) are becoming increasingly widespread. Organizations that want to use AI for productivity now face an important decision. They can subscribe to commercial LLM services or deploy models on their own infrastructure. Cloud services from providers such as OpenAI, Anthropic, and Google are attractive because they provide easy access to state-of-the-art models and are easy to scale. However, concerns about data privacy, the difficulty of switching service providers, and long-term operating costs have driven interest in local deployment of open-source models. This paper presents a cost-benefit analysis framework to help organizations determine when on-premise LLM deployment becomes economically viable compared to commercial subscription services. We consider the hardware requirements, operational expenses, and performance benchmarks of the latest open-source models, including Qwen, Llama, Mistral, and etc. Then we compare the total cost of deploying these models locally with the major cloud providers subscription fee. Our findings provide an estimated breakeven point based on usage levels and performance needs. These results give organizations a practical framework for planning their LLM strategies.
Index Terms:
Large Language Models, On-Premise Deployment, Cost-Benefit Analysis, Total Cost of Ownership.
I Introduction
The rapid advancement of LLMs has encouraged companies and organizations to integrate them into user-facing services, which enables more responsive and adaptive interactions that cater to diverse user needs [1, 3, 2]. As adoption grows, organizations face a critical strategic choice: whether to rely on commercial cloud-based services subscription or to invest in their own on-premise deployment infrastructure [4, 5]. Through APIs and subscription services, providers like OpenAI, Anthropic, and Google are making their state-of-the-art models easy to access. However, the expenses of these solutions can rise rapidly when used by a large group of people [6]. They also bring up issues with regulatory compliance, data protection, and the challenge of transferring service providers [7, 8]. For instance, Desai et al. [9] point out that privacy concerns are a key factor slowing the adoption of LLMs in financial organizations where compliance and trust are crucial.
At the same time, organizations now have the option to deploy competitive open-source models released by the open-source community, including LLaMA [10], Mistral [11], and Qwen [12]. Recent advances in GPU hardware (NVIDIA H100[13], AMD MI300X[14]) and inference optimization frameworks (vLLM[15], NVIDIA TensorRT-LLM[16], DeepSpeed[17]) also contributed to making local deployment more feasible. As a result, it is worth for enterprises to reconsider whether building their own AI infrastructure may be a more economically viable option.
Despite an increasing interest in balancing online subscriptions with local deployment, a lack of systematic, quantitative analysis comparing the two approaches remains. Organizations need a clear framework to evaluate the break-even point when local deployment becomes cost-effective compared to commercial services.
This paper provides a comprehensive cost-benefit analysis framework for on-premise open-source LLM deployment. Our main contributions are:
  1. 1.
    A systematic survey of current commercial LLM pricing models and open-source alternatives suitable for local deployment.
  2. 2.
    Mathematical models for total cost of ownership (TCO) analysis comparing local open-source LLM deployment and commercial API usage.
  3. 3.
    An online playground [^1] Available at: https://v0-ai-cost-calculator.vercel.app/ where enterprise users can apply the cost-benefit framework to the latest models, explore hardware/API trade-offs, and tailor break-even analysis to their workloads.
Our analysis reveals that on-premise deployment are economically viable, with break-even periods typically within a few months for small models, 2 years for medium models and 5 years for larger models, making it viable primarily for organizations with extreme high-volume processing requirements ($\geq$50M tokens/month) or strict data residency mandates.
II Background and Related Work
II-A LLM Deployment Paradigms
Today, modern LLM deployment follows three paradigms: cloud, on-premise, and hybrid. Cloud services provide direct access to state-of-the-art models without pre-setup, but introduce challenges such as recurring costs, data security, and sovereignty [19, 18]. On-premise deployments, in contrast, give organizations full control over infrastructure and ensure that sensitive data remain in-house rather than being used for external model training. For domains such as healthcare, finance, and law, local deployment is often preferred due to strict security and compliance requirements [20]. However, local setups require significant upfront investment and specialized expertise. To balance these trade-offs, hybrid approaches combine local execution for sensitive data with selective cloud offloading for scalability, latency reduction, and accuracy [21].
II-B Cost Analysis Literature
Companies and organizations deploying LLMs typically weigh performance and governance, but cost consistently stands out as a top concern. Prior research into LLM economy has largely focused on four key areas. Firstly, some studies aim to reduce costs without loss of accuracy, concentrating on API pricing and usage tactics like model cascades [6, 22]. Secondly, systems-level effort emphasizes on reducing inference cost per token, achieved through optimizations such as paging-based memory management, quantization, batching, and speculative decoding [15, 23, 24, 25]. Thirdly, deployment-model research examines the financial aspects of serverless inference, provisioned throughput, and adapter-sharing for multi-tenant workloads [26, 27, 28, 29]. Finally, recent researchers have modeled the total cost of ownership (TCO). This work underscores the trade-offs between spending on cloud-based API and amortized hardware costs of on-premises local deployment. They also discussed the impact of precision formats (e.g., FP8) and energy efficiency [30, 31, 32].
Despite this progress, few frameworks truly integrate the economics of cloud and on-premise. While extensive literature already covers individual topics (e.g., cloud API costs or local hardware efficiency), a unified model to guide business decisions is still missing. This study aim to bridge this gap. We seek to embed cost analysis within a comprehensive framework that considers pricing, performance, and governance together.
II-C Open-Source Model Ecosystem
The open-source LLM ecosystem has accelecrated, with many models now achieving performance as well as commercial offerings. Meta’s LLaMA-3 family provides models ranging from 8B to 405B parameters. Outperforming or matching with the commercial giants like Claude and Gemini’s 70B and 405B in reasoning and knowledge tests [33]. Alibaba’s Qwen series reinforces this pattern. Its Qwen-3 model set new standards in mathematics, coding, and multilingual understanding while remaining openly available [34].
The quick spread of powerful models makes them easier to adopt. Organizations can now run advanced LLMs themselves, which provides more control over expenses, tailoring new features, and data security. Consequently, organizations used to depend on commercial APIs are now seriously looking at open-source alternatives for their operation needs.
III Survey Methodology and Model Selection
III-A Performance Evaluation Framework
We’ve developed a thorough performance evaluation framework to ensure our cost-benefit analysis accurately reflects real-world deployment. Our assessment criteria are:
Accuracy Metrics. Selecting the right model begins with accuracy, but accuracy in isolation is insufficient for enterprise decision-making. Organizations require evidence of model performance on tasks directly relevant to their workflows. To select models that are likely to deliver practical value, we ground our evaluation in standardized benchmarks that represent complex analytical, quantitative, and technical challenges. Specifically, we adopt GPQA (graduate-level reasoning)[59], MATH-500 (mathematical problem solving)[60], MMLU-Pro (broad multi-task language understanding)[63], and LiveCodeBench (software engineering and debugging tasks)[61]. Each benchmark captures a distinct class of enterprise need: high-level research and analysis (GPQA), computationally intensive problem solving (MATH-500), cross-domain reasoning (MMLU-Pro), and applied software development support (LiveCodeBench). We then compare accuracy across commercial systems (e.g., OpenAI, Anthropic) and leading open-source models (e.g., LLaMA, QWen-3), quantifying the performance gap and assessing whether open alternatives can feasibly replace commercial APIs. The benchmark results are reported in Table I(the data are adapted from Artificial Analysis [62]), while the performance gap between open-source and commercial models is summarized in Table VI. By tying evaluation to representative tasks and aligning results with enterprise priorities, this framework ensures that model selection is not only about raw accuracy but also about measurable contributions to organizational value, balanced against cost and deployment flexibility.
III-B Model Selection Criteria
Bringing accuracy and operations together, we define the following criteria for models included in our survey:
  1. 1.
    Performance Parity: Benchmark scores within 20% of leading commercial models. This threshold reflects enterprise practice, where modest performance gaps are often outweighed by considerations of cost, security, and integration.
  2. 2.
    Deployment Feasibility: Hardware requirements suitable for typical enterprise environments
  3. 3.
    License Compatibility: Open-weight models distributed under permissive licenses that allow commercial deployment.
  4. 4.
    Community Support: Evidence of an active optimization and development ecosystem, ensuring ongoing improvements and stability.
TABLE I: Performance Benchmarks for Trending Models(data from Artificial Analysis [62])
Model Total Parameters GPQA MATH-500 LiveCodeBench MMLU-Pro
Large Open Models
Kimi-K2[35, 36] 1T 76.6% 97.1% 55.6% 82.4%
GLM-4.5[37, 38] 355B 78.2% 97.9% 73.8% 83.5%
Qwen3-235B[39, 40] 235B 79% 98.4% 78.8% 84.3%
Medium Open Models
gpt-oss-120B[41, 42] 120B 78.2% 63.9% 80.8%
GLM-4.5-Air[37, 43] 106B 73.3% 96.5% 68.4% 81.5%
Llama-3.3-70B[44, 45] 70B 49.8% 77.3% 28.8% 71.3%
Small Open Models
EXAONE 4.0 32B[46, 47] 32b 73.9% 97.7% 74.7% 81.8%
Qwen3-30B[39, 48] 30B 70.7% 97.6% 70.7% 80.5%
Magistral Small[49, 50] 24B 64.1% 96.3% 51.4% 74.6%
Commercial Reference
GPT-5 85.4% 99.4% 66.8% 87.1%
Claude-4-Sonnet 68.3% 93.4% 44.9% 83.7%
Claude-4-Opus 70.1% 94.1% 54.2% 86.0%
Grok-4 87.7% 99.0% 81.9% 86.6%
Gemini-2.5-Pro 84.4% 96.7% 80.1% 86.2%
Table I presents our selected models and their benchmark performance compared to commercial alternatives.
IV Commercial LLM Pricing Models
The API subscription model charges based on the number of tokens processed (input and output), suitable for integrating LLM capabilities into custom applications. Costs can vary dramatically depending on usage patterns, batch processing, and model selection. We provide detailed cost breakdowns and usage scenarios in TableII.
TABLE II: Commercial LLM API Pricing (per 1M tokens, USD)
Provider Model Input Output
OpenAI[51] GPT-5 \$1.25 \$10.00
Anthropic[52] Claude-4 Opus \$15.00 \$75.00
Anthropic Claude-4 Sonnet \$3.00 \$15.00
xAI[53] Grok-4 \$3.00 \$15.00
Google[54] Gemini 2.5 Pro \$1.25 \$10.00
V On-Premise Deployment: A Cost Breakdown
On-premise deployment means running LLMs entirely using an organization’s own data centers or specially designed hardware. This approach does not require any external cloud providers, which provides full control of privacy. We decompose the cost structure into these 3 aspects:
  • Capital Expenditures (CapEx): Hardware (GPUs, servers, storage), initial setup, networking.[55]
  • Operational Expenditures (OpEx): Electricity, cooling, maintenance, personnel, software licensing.[55]
  • Scaling Costs: Additional hardware and operational costs as user base or workload grows.
V-A Capital Expenditures: The Hardware
The largest upfront cost is compute infrastructure. GPU selection is crucial:
  • Data Center Grade (e.g., NVIDIA A100[57]): Highest performance, best for large-scale, multi-user deployments. Enables running the largest models and highest throughput.
  • Prosumer/Workstation Grade (e.g., RTX 5090[56]): Suitable for small teams or research, lower cost but limited scalability.
TABLE III: Comparison of A100-80GB and RTX 5090-32GB GPUs
GPU Memory Power Price (USD)
NVIDIA RTX 5090-32GB[56] 32 GB 575 W \$2,000
NVIDIA A100-80GB[57] 80 GB 400 W \$15,000
Table III lists the memory, power, and approximate price for NVIDIA 5090-32GB and A100-80GB GPUs. Table IV lists the hardware resources needed to set up the service.
TABLE IV: Hardware Requirements for Deployable Open-Source Models (FP8/W8A16 on A100[57]/RTX 5090[56])
Model MoE Total Params Active Params VRAM (FP8) Hardware Deployment Tokens/sec
Large Open Models
[…]
Medium Open Models
[…]
Small Open Models
[…]
V-B Operational Expenditures: The Hidden Costs
Local deployment does not only involve upfront hardware purchases; the running costs are also not negligible. In practice, operational expenditures come from many sources, including network, storage, backup, security, and facility overhead. In this paper, we restrict attention to the electricity cost from model execution, to keep our cost model tractable and transparent. Power usage is the dominant recurring expense and the easiest to quantify.
VI Cost Model and Break-Even Analysis
We developed a quantitative model to assess the cost and feasibility of matching the throughput of leading commercial LLMs. Our model uses available benchmark, hardware, and pricing data.
VI-A Cost-Performance Tradeoffs of Open-Source LLM Deployment
Table I and table VI highlights that open-weight models can deliver competitive performance. While the largest open models (e.g., Kimi-K2, GLM-4.5, Qwen3-235B) require multi-node GPU clusters costing upwards of \$200k, their accuracy on enterprise-relevant benchmarks place them within striking distance of the strongest closed models. This demonstrates that, from a purely technical standpoint, open alternatives can approach commercial state-of-the-art, albeit with greater operational complexity.
More importantly, the gap between “large” and “medium” deployments is narrower than might be expected. Medium-scale models such as gpt-oss-120B, GLM-4.5-Air, and Llama-3.3-70B can run comfortably on just two A100-80GB GPUs (\$30k), with accuracy reduction typically within 10%. The cost of ownership is thus significantly lower while still delivering practical accuracy on reasoning, coding, and domain-specific tasks.
Finally, small models (e.g., EXAONE 4.0 32B, Qwen3-30B, Magistral Small) demonstrate that even sub-30B parameter deployments are feasible on a single consumer-grade RTX 5090 (\$2k). With comparable performance to medium open models, the small models are well-suited for small and medium enterprises that prioritize cost efficiency and local control. While they fall short of the very highest benchmarks, the observed performance differences between 30B- and 70B-class models are relatively modest in practice, suggesting that smaller deployments can already cover a broad range of enterprise use cases.
While commercial APIs retain a slim lead in peak accuracy and efficiency, open-weight models now offer viable, cost-effective alternatives. The performance gap between large, medium, and small open deployments is far narrower than the order-of-magnitude differences in hardware cost. For many organizations, deploying a medium or even small open model locally provides a sustainable “break-even” alternative that balances capability, cost, and independence from external providers.
TABLE V: Hardware and Capacity Analysis for Open-Source Models (Electricity at \$0.15/kWh)
Model Hardware Cost Electricity Cost/Month Throughput (tok/sec) Token Capacity/Month
Large Open Models
Kimi-K2 \$240k \$126.72 800 506.9M
GLM-4.5 \$90k \$47.52 400 253.4M
Qwen3-235B \$60k \$31.68 400 253.4M
Medium Open Models
gpt-oss-120B \$30k \$15.84 220 139.4M
GLM-4.5-Air \$30k \$15.84 200 126.7M
Llama-3.3-70B \$15k \$7.92 190 120.4M
Small Open Models
EXAONE 4.0 32B \$2k \$13.20 200 126.7M
Qwen3-30B \$2k \$13.20 180 114.0M
Magistral Small \$2k \$13.20 150 95.0M
VI-B Cost Model
The initial one-time infrastructure setup cost is defined as:
$C_{\text{hardware}}=N_{\text{GPU}}\cdot C_{\text{GPU}}$ (1)
Based on operational requirements for business hours (8 hours/day, 20 days/month), the monthly electricity cost for LLM inference is defined as:
$C_{\text{electricity}}=N_{\text{GPU}}\cdot P_{\text{GPU}}\cdot H_{\text{operation}}\cdot R_{\text{electricity}}$ (2)
Given the cost of infrastructure setup and monthly electricity cost for inference, we define the local deployment cost as:
$C_{\text{local}}(t)=C_{\text{hardware}}+C_{\text{electricity}}\cdot t$ (3)
where:
  • $t$: Number of months after deployment
  • $C_{\text{GPU}}$: Cost per GPU (e.g., A100-80G or RTX 5090-32G)
  • $N_{\text{GPU}}$: Number of GPUs
  • $P_{\text{GPU}}$: GPU power consumption (Watts)
  • $R_{\text{electricity}}$: Business electricity rate (\$/kWh)
  • $H_{\text{operation}}$: Operating hours per month
  • $T_{\text{throughput}}$: Token generation rate (tokens/second)
VI-C Commercial API Throughput Analysis
To compare against local deployments, we normalize costs by the token generation capacity $Q_{capacity}$ achievable on open-source hardware (from Table IV).
For a commercial API, the equivalent cost of generating the same number of tokens per month is defined as:
$C_{\text{API}}(Q_{\text{capacity}})=\frac{Q_{\text{capacity}}}{3}\cdot\frac{C_{\text{input}}}{1M}+\frac{2Q_{\text{capacity}}}{3}\cdot\frac{C_{\text{output}}}{1M}$ (4)
Given the cost of generating the equivalent number of toekns per month with the open-source models, we define the API usage cost as:
$C_{API}(t)=C_{API}(Q_{capacity})\times t$ (5)
where:
  • $Q_{capacity}$: Monthly token generation capacity of the open model and defined as:
    $Q_{\text{capacity}}=T_{\text{throughput}}\cdot H_{\text{operation}}\cdot 3600$ (6)
  • $C_{input}$: API price for 1M input tokens
  • $C_{output}$: API price for 1M output tokens
  • $2:1$ ratio reflects typical workloads where completions are longer than prompts
Thus, for any open model with capacity $Q_{capacity}$, the equivalent API expenditure scales linearly with both token throughput and the API pricing structure. This allows direct comparison of one-time hardware investment plus electricity costs versus recurring API charges.
VI-D Break-Even Intersection Analysis
We express local and API costs as time-dependent functions to identify the operating horizon where local deployment becomes more economical.
  • $C_{local}(t)$ is a step function starting at hardware cost, growing linearly with electricity costs.
  • $C_{API}(t)$ is a linear function that grows with the usage of token usage capacity.
To achieve break-even condition, we need to find $t$ that satisfies:
$C_{local}(t)=C_{API}(t)$ (7)
Solving this equation yields the critical time $t^{*}$, the point at which the cumulative cost of local deployment equals the cumulative API expenditure. For $t>t^{*}$, on-premise deployment becomes more cost-effective, particularly under sustained high-throughput workloads. In the following section, we present the calculated break-even points $t^{*}$ across different open-source and commercial LLMs to quantify the threshold at which local hosting achieves economic advantage.
Figure 1: Break-even analysis for open-source LLMs: Large (top), Medium (middle), and Small (bottom) models.
VI-E Comprehensive Break-Even Analysis Results
By substituting the values from Table II and Table III into Equation (7), we obtain the results presented in Figure 1. A detailed cost breakdown corresponding to these results is provided in Table V. Our break-even analysis across 54 deployment scenarios shows that the economic viability varies widely. This analysis compares nine open-source models with six commercial API services.
TABLE VI: Break-Even Analysis Summary for All Model-API Combinations (months) with Amortized Performance Differences
Open Model GPT-5 Claude-4 Opus Claude-4 Sonnet Grok-4 Gemini 2.5 Pro Range
Large Open Models
Kimi-K2 69.3 (-6.75%) 8.7 (+1.83%) 44.0 (+5.35%) 44.0 (-10.88%) 63.1 (-8.93%) 8.7-69.3
GLM-4.5 51.5 (-1.32%) 6.5 (+7.25%) 32.8 (+10.78%) 32.8 (-5.45%) 47.0 (-3.50%) 6.5-51.5
Qwen3-235B 34.0 (+0.45%) 4.3 (+9.03%) 21.8 (+12.55%) 21.8 (-3.68%) 31.1 (-1.73%) 4.3-34.0
Medium Open Models
gpt-oss-120B 30.9 (-5.47%) 3.9 (+4.20%) 19.8 (+8.67%) 19.8 (-11.10%) 28.2 (-9.27%) 3.9-30.9
GLM-4.5-Air 34.0 (-4.75%) 4.3 (+3.83%) 21.8 (+7.35%) 21.8 (-8.88%) 31.1 (-6.93%) 4.3-34.0
Llama-3.3-70B 17.8 (-27.88%) 2.3 (-19.30%) 11.4 (-15.78%) 11.4 (-32.00%) 16.2 (-30.05%) 2.3-17.8
Small Open Models
EXAONE 4.0 32B 2.26 (-2.65%) 0.3 (+5.93%) 1.4 (+9.45%) 1.4 (-6.43%) 2.06 (-4.48%) 0.3-2.26
Qwen3-30B 2.5 (-5.25%) 0.3 (+3.38%) 1.6 (+6.90%) 1.6 (-9.00%) 2.3 (-7.05%) 0.3-2.5
Magistral Small 3.0 (-12.25%) 0.4 (-3.23%) 1.9 (+0.28%) 1.9 (-15.73%) 2.76 (-13.78%) 0.4-3.0
Small Model Deployment Economics: Small Model Deployment Economics: Small open-source models are highly cost-effective, often break-even in less than 3 months. For example, the EXAONE-4.0-32B breaks even against Claude-4 Opus in 0.3 months, and against other commercial APIs within a favorable 2-3 months. This allows smaller companies to maintain their competitive AI without the recurring costs.
Medium-Scale Deployment Analysis: Medium-Scale Deployment Analysis: For medium models, break-even periods range from 3.8 to 31.2 months. The GLM-4.5-Air model has the shortest payback against Claude-4 Opus at 3.8 months. Llama-3.3-70B against Gemini 2.5 Pro requires the longest breakeven period at 31.2 months. The \$15,000-\$30,000 hardware cost and \$7.92-\$15.84 monthly operating costs create a significant economic barrier for sustained use.
Large Model Economic Challenges: Large-scale models present the most challenging economic conditions, with break-even horizons stretching anywhere from just over 3 months to nearly 9 years. Among them, the Qwen3-235B illustrates the more favorable end of the spectrum, reaching cost recovery in about 3.5 months when compared against a premium service such as Claude-4 Opus. At the opposite extreme, models like Kimi-K2 illustrate the difficulty of competing with aggressively priced providers: when benchmarked against Gemini 2.5 Pro, the time to break even lengthens dramatically, in this case to over 108 months.
VI-F Commercial API Competitive Analysis
The break-even analysis reveals distinct pricing tiers among commercial providers that fundamentally influence deployment decisions.
Premium Tier - Claude-4 Opus: At the top end, Claude-4 Opus is the most expensive service, averaging \$45 per million tokens. Because of this high cost, local deployment becomes attractive across all model sizes. Break-even can occur in as little as 0.3 months for small models and 3.5–6.9 months for large ones.
Competitive Tier - Claude-4 Sonnet, Grok-4: These providers fall into the middle range, with prices between \$3.13 and \$9.00 per million tokens. Against these baselines, break-even stretches from 1.4 months up to 44.1 months. This exerts moderate economic pressure on medium- and large-scale open-source models, yet it does not eliminate the rationale for local deployment.
Cost-Leadership Tier - Gemini 2.5 Pro, GPT-5: Gemini 2.5 Pro and GPT-5 represents the most cost-competitive commercial option. This aggressive pricing extends break-even periods substantially, ranging from 3.0 months for small models to 63.3 months for large models, challenging the economic viability of on-premise deployment.
VI-G Deployment Decision Framework
We analyzed break-even points by both model size and enterprise type. Small, medium, and large organizations have different needs for computing power, regulatory compliance, and financial resources. These factors shape whether on-premise deployment is practical. Table VI reports the quantitative results. The following sections explain what these results mean for organizations of different sizes.
VI-G1 Small Enterprises (SMEs)
For SMEs with limited budgets and moderate workloads (<<10M tokens/month), small open-source models such as EXAONE 4.0 32B and Qwen3-30B offer the most viable entry point. Our results show that break-even can occur in as little as 0.3–3 months depending on the commercial baseline (e.g., Claude-4 Opus vs. GPT-5). This aligns with recent findings that SMEs prioritize cost savings and data control over absolute performance [20, 22]. The feasibility of deploying on consumer-grade GPUs (e.g., RTX 5090 at $\sim$\$2,000) further reduces capital barriers. Typical use cases include customer support automation, internal knowledge search, and lightweight document analysis—tasks where 30B-class models provide sufficient accuracy [34, 46].
VI-G2 Medium Enterprises
Medium-scale enterprises (processing 10–50M tokens/month) represent the sweet spot for on-premise adoption. Medium models such as GLM-4.5-Air and Llama-3.3-70B demonstrate balanced economics, with break-even periods ranging from 3.8 to 34 months depending on provider comparison. Hardware requirements remain manageable (\$15k–\$30k for dual A100 setups), and throughput levels (120–220 tokens/sec) are sufficient for concurrent business workloads such as code assistance, analytics, and customer-facing applications. This tier benefits most from hybrid strategies, where sensitive workloads run locally and burst traffic leverages cloud APIs [21, 7]. Regulatory-driven industries such as healthcare and finance particularly favor medium-scale deployments for balancing compliance and cost [9].
VI-G3 Large Enterprises
For large enterprises with extreme-scale workloads (>>50M tokens/month), large open-source models (e.g., Qwen3-235B, Kimi-K2) become economically attractive, albeit with longer break-even horizons (3.5–69.3 months). While upfront investments exceed \$40k–\$190k, these organizations often already operate GPU clusters for other workloads, reducing incremental CapEx [55, 30]. Use cases include enterprise-wide generative applications, advanced research, and domain-specific copilots requiring reasoning depth comparable to commercial APIs. However, when compared against aggressively priced providers such as Gemini 2.5 Pro, break-even can extend to 5–9 years, challenging the economic case unless privacy, sovereignty, or vendor lock-in are overriding concerns [18, 32]. Thus, for this tier, non-financial factors (e.g., strategic autonomy, compliance) often weigh more heavily than pure cost.
VII Conclusion
This study looks at 54 deployment scenarios and offers organizations a clearer sense of how they might approach LLM application. The findings suggest the economics are not as straightforward as they first seem, and some of the usual assumptions about the practicality of on-premise deployment may not hold up.
  • Small-scale deployments can achieve break-even within as little as 0.3 months relative to premium commercial services, and within at most 3 months under less favorable conditions. This makes local deployment far more accessible than often assumed.
  • Medium-scale deployments demonstrate break-even periods ranging from 2.3 to 34.0 months, creating viable alternatives for organizations with sustained volume requirements
  • Large-scale deployments face break-even periods extending to 69.3 months against cost-competitive services, limiting viability to exceptional organizational circumstances
  • Commercial provider pricing heterogeneity creates order-of-magnitude variations in deployment economics, with premium services enabling rapid cost recovery for local alternatives
Strategic Decision Framework: The study groups deployment scenarios into three categories: those that can pay off quickly (within 0–6 months), those that are longer-term bets (6–24 months), and those that are hard to justify economically (more than 24 months). This approach helps organizations compare different options by weighing both break-even numbers and broader strategic needs.
Market Evolution Context: The costs and benefits of deploying large language models are changing quickly, driven by better models, more efficient hardware, and shifting commercial prices. Standard benchmarks—like those from Artificial Analysis—only show a momentary snapshot, and can become outdated as soon as new optimizations or models appear. This means that choosing how to deploy LLMs isn’t a one-time decision, but more like a long-term investment that has to adapt as the technology and market evolve.
At the same time, it’s becoming increasingly practical for smaller organizations to run their own models on-site, reflecting a broader trend of making LLM technology more accessible. Yet, aggressive pricing from commercial providers continues to push costs down, creating pressure on those who want to run models locally.
Future Research Directions: Future work should empirically validate break-even projections through longitudinal studies of real-world deployments. Total cost of ownership modeling needs to be expanded to incorporate operational realities such as staffing, energy consumption, hardware failures, and maintenance overheads. Hybrid deployment paradigms deserve further exploration, balancing economic efficiency with reliability and strategic flexibility. Finally, new benchmarking frameworks should be developed to track performance and cost metrics over time, enabling fairer and more accurate comparison across the rapidly evolving open-source and commercial ecosystems.
By situating break-even analyses within the broader context of evolving benchmarks, infrastructure costs, and deployment strategies, this survey highlights both the opportunities and the challenges of navigating the shifting landscape of LLM economics.